4 -3 .[2 (3-5) +1]
My solution was
1. [2 (3 -5)
1. [2 (-2) +1]
2. -2= -4 +1 = -5
The answer was wrong, where did I make mistakes? Thank you

Answer by Alan3354(69443)

4 -3 .[2 (3-5) +1]
My solution was
1. [2 (3 -5) **** It appears you did 4-3 first.
1. [2 (-2) +1]
2. -2= -4 +1 = -5
The answer was wrong, where did I make mistakes? Thank you
=============
4 -3*(2 (3-5) +1) Use * for multiply
= 4 -3*(2(-2) +1)
= 4 -3*(-4 +1)
= 4 -3*(-3)
= 4 + 9
= 13